# Gravitational wave probes of particle dark matter: a review

### v1

@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition. Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed a model of the stochastic GW background arising from isolated BHs formed in two ways: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. The major reason for the inability to robustly exclude boson masses is that there exist degeneracies between the boson mass and BH astrophysical quantities, such as spins. The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $[0.8,6]\times 10^{-13}$ eV at 95% confidence. Additionally, Ref. Yuan et al. (2022) has also disfavored bosons with masses between $[1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors. Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al. (2024a), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al. (2024a).
@@ VI.4 Prospects for GW probes of boson clouds
At the moment, despite the vast theoretical playground of ultralight dark matter boson clouds, only a few searches have actually been performed for GWs due to boson-boson annihilations. This is primarily due to the fact that the rich physics present in boson cloud binary black hole systems, and from transitions in isolated ones, can only be probed in space-based GW detectors, which are expected to be operational in the 2030s. Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches. However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen and East (2020); Siemonsen et al. (2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al. (2023, 2024). Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond. Such known unknowns limit potential constraints to be quite model agnostic (not so astrophysically informative) or too tied to uncertain BH population models. These conundrums are not yet resolved, and probably will not be until we start detecting significantly more BHs than we currently do.
@@
All-sky searches for such boson cloud systems could also be enhanced. At the moment, especially in light of boson self-interactions that could alter the GW signal significantly. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2024), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time. Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal.
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors encourage that posteriors on mass and spin be made publicly available in order to incorporate into their future analyses. Furthermore, Ref. LABEL:Guo:2023lbv shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
@@ IX.1 Background
It was first recognized in Gondolo and Silk (1999) that BHs in the DM halo accreting DM adiabatically (slowly) could allow a high-density region of DM to form, called a “spike”, which could enhance the DM annihilation rate inside the spike. Subsequent studies showed that if such spikes formed around supermassive BHs at the center of galaxies, they would would likely have been disrupted during host galaxy mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005). Though, recently, upper bounds on the DM densities for supermassive BHs have been forecasted with LISA Daniel et al. (2025). However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005), since their spikes may not have been torn apart through mergers over cosmic time. Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2024); Dosopoulou and Silk (2025).
@@ IX.2 Dephasing due to dynamical friction of IMRI systems
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit. Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Ghodla (2024), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
PBHs could also host DM spikes, and must do so, since, at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence DM spikes around binary PBH systems, where the primary has a mass of $[1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, just as in astrophysical binary black hole systems, these exotic mergers would be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. Multi-band observation of inspirals/mergers constraining DM spikes are also possible with future deci-hertz and ground-based GW interferometers, which would enhance the constraints on the power-law index and spike-density factors by orders of magnitude relative to individual constraints from each of these detectors Tahelyani et al. (2024); Chen et al. (2024).
@@
Perhaps more surprising, it has been proposed that equal-mass systems could sustain DM spikes even in the last ten orbits of binary black hole merger Aurrekoetxea et al. (2024b). Initially, it had been shown through N-body simulations that heavy, particle-like DM environments would be disrupted well before equal-mass binary black hole systems merge Bertone and Merritt (2005); Kavanagh et al. (2018). But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments,Koo et al. (2024); Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020) or with boson clouds Cao and Tang (2023); Traykova et al. (2023); Chia et al. (2023); Baumann et al. (2019a, 2020); Ikeda et al. (2021); Cardoso et al. (2020); Hannuksela et al. (2019); Baumann et al. (2022b); Zhang and Yang (2020); Kumar Poddar et al. (2020); Wong (2019, 2020); Kavic et al. (2020); East (2018); Siemonsen and East (2020); Baumann et al. (2022a); Siemonsen et al. (2023b); Tomaselli et al. (2023); Poddar et al. (2024); Cao et al. (2024); Guo et al. (2024b), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al. (2023). Ref. Aurrekoetxea et al. (2024b) found that amount of dephasing that exists within the last ten orbits of a binary black hole merger depends strongly the scalar DM mass and the masses of the BHs:
@@ IX.3 Upper limits on BH environments
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small with respect to the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3 Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Kavanagh et al. (2020).
@@ XI Conclusions
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^7] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.

### v3

@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals. Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed a model of the stochastic GW background arising from isolated BHs formed in two ways: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. The major reason for the inability to robustly exclude boson masses is that there exist degeneracies between the boson mass and BH astrophysical quantities, such as spins. The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $[0.8,6]\times 10^{-13}$ eV at 95% confidence. Additionally, Ref. Yuan et al. (2022) has also disfavored bosons with masses between $[1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors. Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al. (2024b), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al. (2024b).
@@ VI.4 Prospects for GW probes of boson clouds
At the moment, despite the vast theoretical playground of ultralight dark matter boson clouds, only a few searches have actually been performed for GWs due to boson-boson annihilations. This is primarily due to the fact that the rich physics present in boson cloud binary black hole systems, and from transitions in isolated ones, can only be probed in space-based GW detectors, which are expected to be operational in the 2030s. Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches. However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen and East (2020); Siemonsen et al. (2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al. (2023, 2025). Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond. Such known unknowns limit potential constraints to be quite model agnostic (not so astrophysically informative) or too tied to uncertain BH population models. These conundrums are not yet resolved, and probably will not be until we start detecting significantly more BHs than we currently do.
@@
All-sky searches for such boson cloud systems could also be enhanced. At the moment, especially in light of boson self-interactions that could alter the GW signal significantly. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2025), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time. Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal.
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors encourage that posteriors on mass and spin be made publicly available in order to incorporate into their future analyses. Furthermore, Ref. Guo et al. (2024c) shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
@@ VIII.1 Transmuted black holes in binaries
In contrast to TBHs, we note that if DM collects inside neutron stars, a different signature would be present. Depending on what equation-of-state assumed for baryonic or fermionic DM, neutron stars could actually be DM admixed neutron stars, which would have a particular equation-of-state that differs from the (currently unknown) canonical one. Evidence for such an effect could be found in changes to the tidal deformability parameter that could be measured by GW detectors in the future Collier et al. (2022).
@@ IX.1 Background
It was first recognized in Gondolo and Silk (1999) that BHs in the DM halo accreting DM adiabatically (slowly) could allow a high-density region of DM to form, called a “spike”, which could enhance the DM annihilation rate inside the spike. Subsequent studies showed that if such spikes formed around supermassive BHs at the center of galaxies, they would would likely have been disrupted during host galaxy mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005). Though, recently, upper bounds on the DM densities for supermassive BHs have been forecast with LISA Daniel et al. (2025). However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005); Kim et al. (2023), since their spikes may not have been torn apart through mergers over cosmic time. Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2024); Bertone (2024); Dosopoulou and Silk (2025).
@@ IX.2 Dephasing due to dynamical friction of IMRI systems
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al. (2022). Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Rahman et al. (2024); Chowdhuri et al. (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
PBHs could also host DM spikes, and must do so, since, at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence DM spikes around binary PBH systems, where the primary has a mass of $[1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, just as in astrophysical binary black hole systems, these exotic mergers would be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. Multi-band observation of inspirals/mergers constraining DM spikes are also possible with future deci-hertz and ground-based GW interferometers, which would enhance the constraints on the power-law index and spike-density factors by orders of magnitude relative to individual constraints from each of these detectors Tahelyani et al. (2024); Chen et al. (2024b).
@@
Perhaps more surprising, it has been proposed that equal-mass systems could sustain DM spikes even in the last ten orbits of binary black hole merger Aurrekoetxea et al. (2024b). Initially, it had been shown through N-body simulations that heavy, particle-like DM environments would be disrupted well before equal-mass binary black hole systems merge Bertone and Merritt (2005); Kavanagh et al. (2018). But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments Koo et al. (2024); Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020); Bromley et al. (2024) or with boson clouds Cao and Tang (2023); Traykova et al. (2023); Chia et al. (2023); Baumann et al. (2019a, 2020); Ikeda et al. (2021); Cardoso et al. (2020); Hannuksela et al. (2019); Baumann et al. (2022b); Zhang and Yang (2020); Kumar Poddar et al. (2020); Wong (2019, 2020); Kavic et al. (2020); East (2018); Siemonsen and East (2020); Baumann et al. (2022a); Siemonsen et al. (2023b); Tomaselli et al. (2023); Poddar et al. (2024); Cao et al. (2024); Guo et al. (2024d), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al. (2023). Ref. Aurrekoetxea et al. (2024b) found that amount of dephasing that exists within the last ten orbits of a binary black hole merger depends strongly the scalar DM mass and the masses of the BHs:
@@ IX.3 Upper limits on BH environments
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small with respect to the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3 Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al. (2020).
@@ XI Conclusions
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Croon et al. (2023). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^7] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.